# UniCUE: Unified Recognition and Generation Framework for Chinese Cued Speech Video-to-Speech Generation

### v3

@@ 4.1. Experimental Setting
Architecture Details. The CSV2S pathway is entirely built upon the AudioLDM  (Liu et al., 2023), including its VAE encoder-decoder, latent diffusion model, and vocoder components. For CSR, the Transformer in visual process, tokenizer, text-ViT, and text decoder are initialized from MBart (Liu et al., 2020b). Detailed training and inference configurations are provided in Appendix Section 1.
Evaluation Metrics. We evaluate the synthesized speech from three perspectives: linguistic accuracy, temporal synchronization, and speech quality. Linguistic accuracy is quantified by the Word Error Rate (WER) between the recognized text and ground truth. Temporal synchronization is assessed using SyncNet (Chung and Zisserman, 2016), reporting LSE-D (temporal distance) and LSE-C (confidence score). Speech quality is evaluated via STOI (Taal et al., 2010) for intelligibility and DNSMOS (Reddy et al., 2021) for naturalness.
Comparison Methods. We evaluate our UniCUE against: (1) CSV2S (Ours): direct speech synthesis without CSR assistance; (2) CSR (Ours): including pose-aware visual processor, text encoder and decoder, and semantic alignment pool; (3) CSR methods: CMML (Liu and Liu, 2023) and EcoCued (Liu et al., 2024); (4) V2S methods: Lip2Speech (Choi et al., 2023) and LipVoicer (Yemini et al., 2024).
4.2. Comparison with SOTA Methods
@@
Table 3. Ablation Studies of model components on test data of norma hearing cuers and hearing-impaired cuers. The notations X††, X, and X* indicate ablated versions of the architecture X, where the pose maps, semantic alignment pool, and VPA module are removed, respectively.
@@
Figure 5. User study results for accuracy, quality, and synchronization metrics on normal-hearing (a) and hearing-impaired (b) test samples.
4.3. Ablation Studies
@@ 5. Conclusion
This work introduces UniCUE, the first unified framework for directly generating speech from CS videos. By integrating fine-grained visual understanding with diffusion-based speech synthesis, UniCUE produces intelligible speech with precise temporal alignment. Key components including the pose-aware visual processor, semantic alignment pool, and VisioPhonetic Adapter, enable effective knowledge transfer from CS recognition (CSR) to CS video-to-speech generation (CSV2S), enhancing both linguistic accuracy and temporal synchronization. Additionally, we introduce the UniCUE-HI corpus, a new CS dataset featuring both normal-hearing and hearing-impaired cuers. Extensive experiments on this dataset demonstrate that UniCUE state-of-the-art methods across multiple evaluation metrics.
Acknowledgements.
@@
This work was supported by the National Natural Science Foundation of China (No. 62471420), GuangDong Basic and Applied Basic Research Foundation (2025A1515012296), and CCF-Tencent Rhino-Bird Open Research Fund.

### v4

@@ Introduction
Architecture Details. The CSV2S pathway is entirely built upon the AudioLDM  (liu2023audioldm), including its VAE encoder-decoder, latent diffusion model, and vocoder components. For CSR, the Transformer in visual process, tokenizer, text-ViT, and text decoder are initialized from MBart (liu2020multilingual). Detailed training and inference configurations are provided in Appendix Section 1.
Evaluation Metrics. We evaluate the synthesized speech from three perspectives: linguistic accuracy, temporal synchronization, and speech quality. Linguistic accuracy is quantified by the Word Error Rate (WER) between the recognized text and ground truth. Temporal synchronization is assessed using SyncNet (Chung16a), reporting LSE-D (temporal distance) and LSE-C (confidence score). Speech quality is evaluated via STOI (taal2010short) for intelligibility and DNSMOS (reddy2021dnsmos) for naturalness.
Comparison Methods. We evaluate our UniCUE against: (1) CSV2S (Ours): direct speech synthesis without CSR assistance; (2) CSR (Ours): including pose-aware visual processor, text encoder and decoder, and semantic alignment pool; (3) CSR methods: CMML (liu2023cross) and EcoCued (liu2024computation); (4) V2S methods: Lip2Speech (choi2023intelligible) and LipVoicer (yeminilipvoicer).
Comparison with SOTA Methods
@@
Table 3: Ablation Studies of model components on test data of norma hearing cuers and hearing-impaired cuers. The notations X††, X, and X* indicate ablated versions of the architecture X, where the pose maps, semantic alignment pool, and VPA module are removed, respectively.
@@
Figure 5: User study results for accuracy, quality, and synchronization metrics on normal-hearing (a) and hearing-impaired (b) test samples.
Ablation Studies
@@ Conclusion
This work introduces UniCUE, the first unified framework for directly generating speech from CS videos. By integrating fine-grained visual understanding with diffusion-based speech synthesis, UniCUE produces intelligible speech with precise temporal alignment. Key components including the pose-aware visual processor, semantic alignment pool, and VisioPhonetic Adapter, enable effective knowledge transfer from CS recognition (CSR) to CS video-to-speech generation (CSV2S), enhancing both linguistic accuracy and temporal synchronization. Additionally, we introduce the UniCUE-HI corpus, a new CS dataset featuring both normal-hearing and hearing-impaired cuers. Extensive experiments on this dataset demonstrate that UniCUE outperforms SOTA methods across multiple evaluation metrics.
Acknowledgments
@@
This work was supported by the National Natural Science Foundation of China (No. 62471420), GuangDong Basic and Applied Basic Research Foundation (2025A1515012296), and 2025 Tencent AI Lab Rhino-Bird Program.
Appendix
Figure 6: Comparison of CSV2S generation at inference stage. (a) The combined CS speech generation pipeline of CSR and TTS. (b) Our UniCUE. During inference, the extracted visual mixture embedding serves as a condition to guide direct speech generation.
1. Details of Training and Inference
Training Details. The training process is divided into two sequential phases to ensure effective knowledge transfer from visual understanding to speech synthesis. Stage 1: CSR focuses on learning fine-grained spatiotemporal representations by jointly optimizing the pose-aware visual processor, the semantic alignment pool, text encoder and decoder. This stage trains the model using the hybrid cross-entropy loss $\mathcal{L}_{R}$ (denoted in main text Equation (1)) and the contrastive alignment loss $\mathcal{L}_{S}$ (denoted in main text Equation (5)), formulated as: $\mathcal{L}=\mathcal{L}_{R}+\mathcal{L}_{S}.$ Stage 2: CSV2S leverages the learned visual representations from Stage 1 to train the VPA network and diffusion model with the objective $\mathcal{L}_{G}$ (denoted in main text Equation (2)).
Inference Details. During inference, the framework operates in a streamlined single-path mode to synthesize speech directly from cued CS videos. As illustrated in Figure 6 (b), given an CS video and its corresponding pose, the trained pose-aware visual processor extracts the visual mixture representation $Z_{mv}$, which encodes both semantic content and cuer characteristics. A noise sample $Z^{t}$, sampled from a standard Gaussian distribution $N(0,I)$, is fed into the trained LDM to predict the speech latent code $\hat{Z_{s}}$ conditioned on the visual mixture representation over $t$ timesteps:
$\hat{Z_{s}}:=\frac{1}{\alpha_{t}}\left(Z_{t}-\left(1-\alpha_{t}\right)\cdot\mathcal{M}(Z^{t},Z_{mv},t)\right),$ (7)
where $\mathcal{M}$ denotes the noise prediction network of the LDM, and $\alpha_{t}$ represents the scheduled variance coefficient at timestep $t$. Subsequently, the predicted speech latent code $\hat{Z_{s}}$ is decoded into a speech mel-spectrogram via VAE decoder, and a pre-trained Vocoder (liu2023audioldm) is used to convert the mel-spectrogram into a speech waveform. Through this series of steps, the CSV2S system is able to generate natural speech that is consistent with the cuer’s characteristics based on the input CS video and its corresponding pose information.
Implementation Details. Raw video frames and pose images are resized to $256\times 256$, while the audio is resampled to 16 kHz and converted into mel-spectrograms. The entire learning framework is implemented in PyTorch, and all experiments are conducted on an NVIDIA RTX A6000. We use the Adam optimizer with a learning rate of $5\times 10^{-4}$ for training, setting the mini-batch size to 2.
Figure 7: Statistics of sentence categories within our UniCue-HI Corpus.
Figure 8: Statistical distributions of samples in the UniCUE-HI corpus.(a) Character-level sentence lengths; (b) Phoneme distribution. Normal-hearing and hearing-impaired groups are distinguished by color. Please zoom in for better view.
2. Details of the UniCUE-HI Corpus
To evaluate model effectiveness on hearing-impaired individuals, we construct the first large-scale Chinese Cued Speech (CS) dataset, UniCUE-HI Corpus, that includes both hearing-impaired and normal-hearing cuers. The proposed UniCUE-HI Corpus is designed to ensure linguistic and contextual diversity, covering a broad range of sentence categories, including daily communication scenarios, as illustrated in Figure 7.
To better understand linguistic variations across user groups, we analyze sentence structure statistics in our dataset. Figure 8 presents the distributions of sentence lengths and phoneme occurrences for normal-hearing and hearing-impaired cuers. While the overall distributions are similar, hearing-impaired users tend to produce slightly shorter sentences on average, indicating potential articulatory constraints and reduced fluency. These characteristics introduce additional modeling challenges and emphasize the importance of including hearing-impaired data for more realistic evaluation.
The dataset was constructed by recruiting 10 cuers (2 normal-hearing and 8 hearing-impaired), each instructed to simultaneously vocalize and encode text sentences sourced from (lei2024bridge). All participants underwent systematic training to ensure accurate and fluent Mandarin Chinese CS production. In addition, we incorporate CS video data from 4 normal-hearing cuers in the MCCS dataset (lei2024bridge), resulting in the UniCUE-HI Corpus with 14 cuers.
All data were collected with informed consent and are publicly released for open-source research purposes.
Figure 9: Visualization of the Mel-spectrograms of the generated speech. Our UniCUE model exhibits superior linguistic accuracy and temporal synchronization compared to baseline methods.
3. Details of Comparison with CSR
Since CSR methods output phoneme-level transcriptions, we design a two-stage pipeline for speech generation, as shown in Figure 6(b). First, phoneme sequences are converted into Chinese characters using a rule-based CS mapping system, which follows a curated phoneme-to-character dictionary and handles tone-based disambiguation. To improve robustness and address homophonic ambiguities, the mapped text is further refined by the QWen language model(qwen). Then, the resulting character sequence is passed into F5-TTS (chen2024f5), a controllable speech synthesizer that generates waveforms conditioned on reference speech from the same speaker. This design ensures a fair comparison with our direct CSV2S model in terms of speaker consistency and prosodic naturalness.
Table 4: Ablation analysis the impact of hand cues in cued speech. (w/o) denotes using lip cues only, without hand gestures. Bold and underlined results are the best and second-best results. $\uparrow$ indicates that larger values are better, while $\downarrow$ indicates that smaller values are preferable.
  Method Normal-hearing cuers Hearing-impaired cuers
[…]
  GT - 7.274 7.314 2.79 - - - - -
CSR(w/o) 0.547 4.692 10.064 1.93 0.29 0.713 3.215 10.032 1.01
CSR 0.186 4.874 9.125 2.53 0.57 0.224 3.342 9.315 2.29
CSV2S (w/o) 0.729 4.614 9.174 1.16 0.09 0.925 3.701 10.124 0.05
CSV2S 0.374 6.245 7.962 2.27 0.42 0.422 5.938 8.347 2.04
UniCUE (w/o) 0.592 4.934 8.927 1.88 0.26 0.793 4.412 9.742 0.09
UniCUE 0.205 6.729 7.632 2.46 0.53 0.248 6.491 8.076 2.17
4. Qualitative Comparison with SOTA Methods
To qualitatively assess the effectiveness of our proposed framework, we visualize the Mel-spectrograms of the generated speech and compare them against those produced by several state-of-the-art (SOTA) methods. As shown in Figure 9, UniCUE generates spectrograms that are more consistent with the ground truth in both temporal structure and phonetic detail. In particular, the generated patterns exhibit clearer formant trajectories and better alignment with visual speech cues, indicating superior temporal synchronization and articulation precision. These visual differences further support the quantitative gains observed in WER and LSE metrics, highlighting UniCUE’s ability to generate intelligible and temporally coherent speech from cued videos.
More generation results are provided in the supplementary demo file.
5. Discussion and Analysis
Visualization Analysis of Semantic Alignment Pool. To assess the effectiveness of the proposed Semantic Alignment Pool (SAP), we present t-SNE visualizations of the learned embeddings across video, pose, and text modalities in Figure 10. Without SAP (Figure 10a), the embeddings are loosely distributed with poor cross-modal alignment. In contrast, enabling SAP (Figure 10b) results in compact and well-aligned clusters across modalities, indicating enhanced semantic coherence. These results demonstrate that SAP effectively bridges the visual-textual modality gap, facilitating grounded multimodal representation learning.
Impact of Hand Cues. Table 4 presents an ablation study assessing the effect of hand cues in CS speech synthesis. Removing hand gestures leads to substantial performance degradation across all metrics, particularly for hearing-impaired cuers. For instance, the WER of CSR increases drastically from 0.224 to 0.713 without hand cues, highlighting their crucial role in compensating for atypical articulatory patterns. Moreover, UniCUE achieves the best results in both LSE-C (6.729 for normal-hearing, 6.491 for hearing-impaired) and STOI (0.53), demonstrating its strong capability in generating expressive and intelligible speech by leveraging multi-modal visual information.
Analysis of Cuer Representation. To quantitatively evaluate the model’s ability to preserve cuer-specific acoustic characteristics, we extract latent cuer embeddings from the generated speech using Resemblyzer (resemblyzer2020). These embeddings are projected into a 2D space via t-SNE (maaten2008visualizing) for visualization. As illustrated in Figure 11, speech samples from the same cuer (e.g., cuers 1 through 6) form compact clusters, while the distances between clusters corresponding to different cuers are notably larger. This demonstrates that the model effectively captures and differentiates personalized acoustic features unique to each cuer.
Figure 10: t-SNE visualization of video, pose, and text embeddings. (a): without Semantic Alignment Pool (SAP). (b): with SAP. Each color represents a different class, and different shapes denote different modalities. The results show that SAP enhances cross-modal alignment by bringing embeddings of the same class closer together in the semantic space.
Figure 11: The t-SNE visualization of 6 normal-hearing cuers embeddings from generated speech. Each point represents a speech sample. Better view by zooming in.
Figure 12: The learning curves for the training and validation stages of our UniCUE model and the single CSV2S model. It can be seen that the efficiency and effectiveness of the unified approach. Better view by zooming in.
Analysis of Computational Efficiency. To quantify the efficiency gains of our unified framework, we evaluate the time costs during both training and inference stages.
(a) Training Efficiency. We compare the training performance of CSV2S and UniCUE. As shown in Figure 12, UniCUE’s training loss converges faster to a stable value within fewer epochs, indicating that the incorporation of fine-grained visual understanding accelerates the learning process. Concurrently, the validation loss exhibits reduced fluctuation, reflecting enhanced model robustness.
(b) Real-time Inference. To assess real-time capability, we compare the inference efficiency of the traditional pipeline (CSR + TTS) with our unified CSV2S framework. UniCUE achieves a significant reduction in inference time (4.15s), representing a 40% speedup over the sequential CSR + TTS approach (6.93s), thereby demonstrating the efficiency benefits of the unified architecture.
Table 5: Ablation results of loss design for CSR. $\mathcal{L}_{\text{CE}}^{\text{masked}}$ denotes masked language modeling loss in CSR objective function. H and HI denote the normal-hearing data and hearing-impaired data, respectively.
[…]
CSR (w/o $\mathcal{L}_{\text{CE}}^{\text{masked}}$) 0.214 0.263
CSR 0.186 0.224
6. Ablation Study of CSR Loss
Table 5 presents an ablation analysis of the hybrid loss design in our CSR module. Removing the masked language modeling loss $\mathcal{L}_{\text{CE}}^{\text{masked}}$ leads to noticeable performance degradation, with WER increasing from 0.186 to 0.214 on normal-hearing data and from 0.224 to 0.263 on hearing-impaired data. This highlights the effectiveness of combining token-level masked supervision with sequence-level training, which enhances contextual understanding and improves transcription accuracy.